# Non-commutative rings with infinitely many maximal subrings

### v1

@@ 3. Rings with Infinitely Many Maximal Subrings
$(1)$ Let $M$ be a maximal left/right ideal of $T$ which is not an ideal of $T$. Since $T$ has only finitely many maximal subring, by $(2)$ of Proposition 3.10, we conclude that $\mathbb{I}(M)/M$ is a finite field. Hence $Char(\mathbb{I}(M)/M)=p$, where $p$ is a prime number. Thus $p.1\in M$, for $1\in\mathbb{I}(M)$. Therefore $pT\subseteq M$.
$(2)$ The proof is similar to $(1)$ and using $(4)$ of Proposition 3.10.
$(3)$ Let $P$ be a left primitive ideal of $T$, which is not a left/right maximal ideal of $T$. Hence there exists a maximal left ideal $M$ of $T$ such that $P=l.ann_{T}(T/M)$. If $M$ is an ideal of $T$, then $M=l.ann_{T}(T/M)=P$, which is absurd. Thus $M$ is not an ideal of $T$ and therefore by $(1)$, we deduce that there exists a prime number $p$ such that $p(T/M)=0$. Hence $p\in P$ and thus $Char(T/P)=p$. ∎
By the previous result, if $T$ is a ring with zero characteristic and there exists a maximal left/right ideal $M$ of $T$ which is not an ideal of $T$ and either $M\cap\mathbb{Z}=0$ or $T/M$ is a torsionfree/faithful $\mathbb{Z}$-module, then $T$ has infinitely many maximal subrings.
@@
Let $T$ be an infinite ring with nonzero characteristic. If $Max_{F}(T):=\{M\in Max(T)\ |\ T/M\ \text{is}\ \text{finite}\}$ is infinite, then $T$ has infinitely many maximal subrings.
@@
Since $n$ has only finitely many prime divisors and for each $M\in Max_{F}(T)$, $Char(T/M)$ is a prime divisor of $n$, we conclude that the exists a prime divisor $p$ of $n$ such that $Max_{p,F}(T):=\{M\in Max_{F}(T)\ |\ Char(T/M)=p\}$ is infinite. Also note that for each $M\in Max_{p,F}(T)$, the ring $T/M$ is a finite simple ring, hence we deduce that there exist a natural number $m$ and a finite field $E$ such that $T/M\cong\mathbb{M}_{m}(E)$. Now we have two cases:
$(1)$ there exist distinct $M_{1},M_{2},\ldots\in Max_{p,F}(T)$ such that $T/M_{i}\cong\mathbb{M}_{m_{i}}(E_{i})$, where $m_{i}>1$ and $E_{i}$ is a field. Thus in this case for each $i$, $T$ has a maximal subring $R_{i}$ which contains $M_{i}$, for $\mathbb{M}_{m_{i}}(E_{i})$ has a maximal subring by $(4)$ of Theorem 2.7. Since for $i\neq j$, $M_{i}+M_{j}=T$, we deduce that $R_{i}\neq R_{j}$. Hence $R_{1},R_{2},\ldots$ are distinct maximal subrings of $T$ and hence we are done.
$(2)$ thus assume that there exist infinitely many distinct maximal ideals $M_{1},M_{2},\ldots\in Max_{p,F}(T)$ such that $T/M_{i}=E_{i}$ is a finite field. Clearly by our assumption $\mathbb{Z}_{p}\subseteq E_{i}$, for each $i\geq 1$. Now if for infinitely many $i$, $\mathbb{Z}_{p}\subsetneq E_{i}$, then similar to proof of $(1)$ we obtain that $T$ has infinitely many maximal subrings (note that, $E_{i}$ is a finite field and $\mathbb{Z}_{p}\subsetneq E_{i}$, therefore $E_{i}$ has a maximal subring, see the comments preceding Theorem 2.7, and hence $T$ has a maximal subring $R_{i}$ which contains $M_{i}$). Thus without lose of the generality we may assume that for each $i\geq 1$, we have $\mathbb{Z}_{p}=E_{i}$. Let $P_{i}=M_{2i}$ and $Q_{i}=M_{2i-1}$, where $i\geq 1$. Then clearly for each $i\geq 1$, we have $T/(P_{i}\cap Q_{i})\cong\mathbb{Z}_{p}\times\mathbb{Z}_{p}$. Since $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ has a maximal subring (in fact, $\{(a,a)\ |\ a\in\mathbb{Z}_{p}\}$ is the only proper and therefore maximal subring of $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, see the comments preceding Theorem 2.7), we conclude that $T$ has a maximal subring $R_{i}$ which contains $P_{i}\cap Q_{i}$. Clearly, whenever $i\neq j$, then $(P_{i}\cap Q_{i})+(P_{j}\cap Q_{j})=T$. Thus we deduce that for $i\neq j$ we have $R_{i}\neq R_{j}$ and therefore $T$ has infinitely many maximal subrings. ∎
It is not hard to see that by the proof of the previous result, in fact there exists a prime divisor $p$ of $n$ such that $|Max_{p,F}(T)|=|Max_{F}(T)|\leq|RgMax(T)|$ for some prime number $p$. Now the next result is in order.
@@
Let $T$ be a ring. If $|Max_{F}(T)|>\aleph_{0}$, then $T$ has infinitely many maximal subrings. In fact, $|Max_{F}(T)|\leq|RgMax(T)|$.
@@
Since $|Max_{F}(T)|>\aleph_{0}$ and the number of prime numbers is countable, we conclude that there exists a prime number $p$ such that $|Max_{p,F}(T)|=|Max_{F}(T)|$. Now by a similar proof of Proposition 3.19, we can see that $|Max_{p,F}(T)|\leq|RgMax(T)|$ and hence we are done. ∎
@@
    If $E_{1}=End_{T}(S_{1})$ is a field which is not an absolutely algebraic field.
@@
    If $E_{1}=End_{T}(S_{1})$ and $E_{2}=End_{T}(S_{2})$ are infinite isomorphic ring.
@@
First note that since $S_{i}$ is a simple left $T$-module, we infer that there exists maximal left ideals $M_{i}$, $i=1,2$, such that $S_{i}\cong T/M_{i}$, as left $T$-module.
$(1)$ Note that $E_{1}^{op}\cong\mathbb{I}(M_{1})/M_{1}$. Now we have two cases, either $M_{1}$ is an ideal of $T$ or not. If $M_{1}$ is an ideal of $T$, then $T/M_{1}\cong E_{1}^{op}$ as ring and therefore $T/M_{1}$ has infinitely many maximal subrings, by [10, Corollary 1.5]. Hence assume that $M_{1}$ is not an ideal of $T$. Since $E_{1}$ is not an absolutely algebraic field, we conclude that $E_{1}$ is an infinite ring. Therefore $\mathbb{I}(M_{1})/M_{1}$ is an infinite ring. Thus by Theorem 3.1, we conclude that $M_{1}]$ and therefore $Max_{l}(T)\setminus Max(T)$ are infinite and hence we are done by $(1)$ of Proposition 3.10.
$(2)$ We have two cases, if $M_{1}$ or $M_{2}$ is not an ideal of $T$. Then similar to last part of the proof of $(1)$ we infer that $T$ has infinitely many maximal subrings. Hence assume that $M_{1}$ and $M_{2}$ are ideals of $T$. In this case we have the following ring isomorphisms
@@
Thus $T/(M_{1}\cap M_{2})\cong E_{1}^{op}\times E_{1}^{op}$, as ring (note $M_{1}\neq M_{2}$, for $S_{1}\ncong S_{2}$). Now since $E_{1}$ and $E_{2}$ are infinite division rings, by Theorem 3.6, we deduce that $T/(M_{1}\cap M_{2})$ has infinitely many maximal subrings. Therefore $T$ has infinitely many maximal subrings. ∎
@@ 4. Rings integral over their centers
In this section we investigate about the number of maximal subrings in a ring $T$ which is integral over its center. First we need some observation and notation. Let $T$ be a ring, then it is clear that $Max_{l}(T)\cap Max(T)=Max_{r}(T)\cap Max(T)=Max_{r}(T)\cap Max_{l}(T)$, we denote this set by $Max_{lr}(T)$ and we define $J^{\prime}(T):=\bigcap_{M\in Max_{lr}(T)}M$. It is clear that $J^{\prime}(T)$ is an ideal of $T$ and $J(T)\subseteq J^{\prime}(T)$. In particular, if $T$ is either left or right quasi duo ring, then $J(T)=J^{\prime}(T)$. Moreover if $Max_{l}(T)\setminus Max(T)=\{M_{1},\ldots,M_{n}\}$ is finite (in particular, by $(5)$ of Proposition 3.10, where $T$ has only finitely many maximal subrings), then $J(T)=J^{\prime}(T)\cap M_{1}\cap\cdots\cap M_{n}$. Note that by Lemma 3.13, if $T$ is an infinite ring with $J^{\prime}(T)=0$, then either $T$ has infinitely many maximal subrings or $C(T)$ is integrally closed in $T$. Now the following is in order.
@@
    $|Max_{lr}(T)|\leq 2^{\aleph_{0}}$. For each $M\in Max_{lr}(T)$, $T/M$ is an absolutely algebraic field (thus $|T/M|\leq\aleph_{0}$).
@@
    for each $M\in Max_{l}^{\prime}(T)\cup Max_{r}^{\prime}(T)$, $J^{\prime}(T)+M=T$.
@@
Assume that $T$ has only finitely many maximal subrings. Thus for each $M\in Max_{lr}(T)$, the division ring $T/M$ has only finitely many maximal subrings. Since $T$ is integral over its center, we conclude that for each $M\in Max_{lr}(T)$, the division ring $T/M$ is integral over its center. Thus by Proposition 3.1, we conclude that $T/M$ is a field. Clearly, $T/J^{\prime}(T)$ embeds in $\prod_{M\in Max_{lr}(T)}T/M$, thus $T/J^{\prime}(T)$ is a commutative reduced ring and therefore $N(T)\subseteq J^{\prime}(T)$. Since $T/J^{\prime}(T)$ is a commutative ring with only finitely many maximal subrings, by [10, Corollary 1.9], we deduce that $T/J^{\prime}(T)$ is a Hilbert ring. By [10, Corollary 1.5], for each $M\in Max_{lr}(T)$, the field $T/M$ is an absolutely algebraic field and therefore $|T/M|\leq\aleph_{0}$. Since $T/J^{\prime}(T)$ is a commutative ring, the first part of $(2)$ is an immediate consequence of [10, (1) of Proposition 2.8]. Therefore we deduce that
$|\frac{T}{J^{\prime}(T)}|\leq|\prod_{M\in Max_{lr}(T)}T/M|\leq 2^{2^{\aleph_{0}}}.$
Hence $(1)$ and $(2)$ hold. Finally for $(3)$, let $M$ be a maximal left ideal of $T$ which is not an ideal of $T$. Since $T/J^{\prime}(T)$ is a commutative ring, we conclude that $J^{\prime}(T)\nsubseteq M$ and therefore $M+J^{\prime}(T)=T$ by maximality of $M$. ∎
As we mentioned in the proof of the previous proposition, if $R$ is a commutative ring with $|Max(R)|>2^{\aleph_{0}}$, then $|RgMax(R)|\geq 2^{\aleph_{0}}$, see [10, (1) of Proposition 2.8]. For non-commutative rings we have the following result.
@@
We prove $(1)$ and $(3)$, the proof of $(2)$ is similar to $(1)$. For $(1)$, first note that if $Max_{l}(T)\setminus Max(T)$ is infinite, then by $(1)$ of Proposition 3.10, we deduce that $T$ has infinitely many maximal subrings. Hence assume that $Max_{l}(T)\setminus Max(T)$ is finite. Thus $|Max_{l}(T)\cap Max(T)|=|Max_{l}(T)|>2^{\aleph_{0}}$. Now since $Max_{lr}(T)=Max_{l}(T)\cap Max(T)$, we deduce that $T$ has infinitely many maximal subrings by $(2)$ of Proposition 4.1.
$(3)$ If $Max(T)\setminus Max_{l}(T)$ is infinite, then by $(5)$ of Proposition 3.10, we conclude that $T$ has infinitely many maximal subrings. Hence assume that $Max(T)\setminus Max_{l}(T)$ is finite and therefore $|Max(T)\cap Max_{l}(T)|=|Max(T)|>2^{\aleph_{0}}$. By a similar proof of $(1)$ we obtain that $T$ has infinitely many maximal subrings. ∎
As the following main result shows if we assume that a ring $T$ is an algebraic $K$-algebra (and therefore is integral over its center), then either $T$ has infinitely many maximal subrings or $T$ has similar algebraic properties like commutative rings with only finitely many maximal subrings, see [10].
@@
    $T$ is a quasi duo ring. In particular, $J(T)=\bigcap_{M\in Max(T)}M$.
@@
    For each $M\in Max(T)$, the ring $T/M$ is an absolutely algebraic field which contains a copy of $K$. In particular, $K$ is an absolutely algebraic field and $|K|\leq|T/M|\leq\aleph_{0}$. Moreover, if $D$ is a division ring which is a subring of $T$, then $D$ is an absolutely algebraic field.
@@
    $T/J(T)$ is a reduced Hilbert commutative ring. In fact, $T/J(T)$ is a von Neumann regular ring which is integral over its prime subring (i.e., $\mathbb{Z}_{p}$, where $p=Char(K)$ is a prime number). Moreover, $T$ is integral over $\mathbb{Z}_{p}$.

### v2

@@ 3. Rings with Infinitely Many Maximal Subrings
$(1)$ Let $M$ be a maximal left/right ideal of $T$ which is not an ideal of $T$. Since $T$ has only finitely many maximal subring, by Proposition 3.10(2)(2), we conclude that $\mathbb{I}(M)/M$ is a finite field. Hence $\operatorname{Char}(\mathbb{I}(M)/M)=p$, where $p$ is a prime number. Thus $p.1\in M$, because $1\in\mathbb{I}(M)$. Therefore $pT\subseteq M$.
$(2)$ The proof is similar to $(1)$ and by the use of Proposition 3.10(4)(4).
$(3)$ Let $P$ be a left primitive ideal of $T$, which is not a left/right maximal ideal of $T$. Hence there exists a maximal left ideal $M$ of $T$ such that $P=\operatorname{l.ann}_{T}(T/M)$. If $M$ is an ideal of $T$, then $M=\operatorname{l.ann}_{T}(T/M)=P$, which is absurd. Thus $M$ is not an ideal of $T$ and therefore by $(1)$, we deduce that there exists a prime number $p$ such that $p(T/M)=0$. Hence $p\in P$ and thus $\operatorname{Char}(T/P)=p$. ∎
By the previous result, if $T$ is a ring of characteristic zero and there exists a maximal left/right ideal $M$ of $T$ that is not an ideal and moreover either $M\cap\mathbb{Z}=0$ or $T/M$ is a torsion-free/faithful $\mathbb{Z}$-module, then $T$ has infinitely many maximal subrings.
@@
Let $T$ be an infinite ring with nonzero characteristic. If $\operatorname{Max}_{F}(T):=\{M\in\operatorname{Max}(T)\ |\ T/M\ \text{is}\ \text{finite}\}$ is infinite, then $T$ has infinitely many maximal subrings.
@@
Since $n$ has only finitely many prime divisors and for each $M\in\operatorname{Max}_{F}(T)$, $\operatorname{Char}(T/M)$ is a prime divisor of $n$, we conclude that the exists a prime divisor $p$ of $n$ such that $\operatorname{Max}_{p,F}(T):=\{M\in\operatorname{Max}_{F}(T)\ |\ \operatorname{Char}(T/M)=p\}$ is infinite. Also note that for each $M\in\operatorname{Max}_{p,F}(T)$, the ring $T/M$ is a finite simple ring, hence we deduce that there exist a natural number $m$ and a finite field $E$ such that $T/M\cong\mathbb{M}_{m}(E)$. Now we have two cases:
$(\bf 1)$ There exist distinct $M_{1},M_{2},\ldots\in\operatorname{Max}_{p,F}(T)$ such that $T/M_{i}\cong\mathbb{M}_{m_{i}}(E_{i})$, where $m_{i}>1$ and $E_{i}$ is a field. Thus in this case for each $i$, $T$ has a maximal subring $R_{i}$ which contains $M_{i}$, for $\mathbb{M}_{m_{i}}(E_{i})$ has a maximal subring by Theorem 2.7(4)(4). Since for $i\neq j$, $M_{i}+M_{j}=T$, we deduce that $R_{i}\neq R_{j}$. Hence $R_{1},R_{2},\ldots$ are distinct maximal subrings of $T$, as desired.
$(\bf 2)$ Thus assume that there exist infinitely many distinct maximal ideals $M_{1},M_{2},\ldots\in\operatorname{Max}_{p,F}(T)$ such that $T/M_{i}=E_{i}$ is a finite field. Clearly by our assumption $\mathbb{Z}_{p}\subseteq E_{i}$, for each $i\geq 1$. Now if for infinitely many $i$, $\mathbb{Z}_{p}\subsetneq E_{i}$, then similar to proof of $(1)$ we obtain that $T$ has infinitely many maximal subrings (note that, $E_{i}$ is a finite field and $\mathbb{Z}_{p}\subsetneq E_{i}$, therefore $E_{i}$ has a maximal subring, see the comments preceding Theorem 2.7, and hence $T$ has a maximal subring $R_{i}$ which contains $M_{i}$). Thus without loss of generality, we may assume that for each $i\geq 1$, we have $\mathbb{Z}_{p}=E_{i}$. Let $P_{i}=M_{2i}$ and $Q_{i}=M_{2i-1}$, where $i\geq 1$. Then clearly for each $i\geq 1$, we have $T/(P_{i}\cap Q_{i})\cong\mathbb{Z}_{p}\times\mathbb{Z}_{p}$. Since $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ has a maximal subring (in fact, $\{(a,a)\ |\ a\in\mathbb{Z}_{p}\}$ is the only proper and therefore maximal subring of $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$, see the comments preceding Theorem 2.7), we conclude that $T$ has a maximal subring $R_{i}$ which contains $P_{i}\cap Q_{i}$. Clearly, whenever $i\neq j$, then $(P_{i}\cap Q_{i})+(P_{j}\cap Q_{j})=T$. Thus we deduce that for $i\neq j$ we have $R_{i}\neq R_{j}$ and therefore $T$ has infinitely many maximal subrings. ∎
It is not hard to see that by the proof of the previous result, in fact there exists a prime divisor $p$ of $n$ such that $|\operatorname{Max}_{p,F}(T)|=|\operatorname{Max}_{F}(T)|\leq|\operatorname{RgMax}(T)|$ for some prime number $p$. Now the next result is in order.
@@
Let $T$ be a ring. If $|\operatorname{Max}_{F}(T)|>\aleph_{0}$, then $T$ has infinitely many maximal subrings. In fact, $|\operatorname{Max}_{F}(T)|\leq|\operatorname{RgMax}(T)|$.
@@
Since $|\operatorname{Max}_{F}(T)|>\aleph_{0}$ and the number of prime numbers is countable, we conclude that there exists a prime number $p$ such that $|\operatorname{Max}_{p,F}(T)|=|\operatorname{Max}_{F}(T)|$. Now by a similar proof of Proposition 3.17, we can see that $|\operatorname{Max}_{p,F}(T)|\leq|\operatorname{RgMax}(T)|$ and hence we are done. ∎
@@
    If $E_{1}=\operatorname{End}_{T}(S_{1})$ is a field which is not an absolutely algebraic field.
@@
    If $E_{1}=\operatorname{End}_{T}(S_{1})$ and $E_{2}=\operatorname{End}_{T}(S_{2})$ are infinite isomorphic ring.
@@
First note that since $S_{1}$ and $S_{2}$ are simple left $T$-modules, we infer that there exist maximal left ideals $M_{1}$ and $M_{2}$, such that $S_{i}\cong T/M_{i}$, as left $T$-modules for $i=1,2$.
$(\bf 1)$ Note that $E_{1}^{op}\cong\mathbb{I}(M_{1})/M_{1}$. Now we have two cases, either $M_{1}$ is an ideal of $T$ or is not. If $M_{1}$ is an ideal of $T$, then $T/M_{1}\cong E_{1}^{op}$ as rings and therefore $T/M_{1}$ has infinitely many maximal subrings, by [10, Corollary 1.5]. Hence assume that $M_{1}$ is not an ideal of $T$. Since $E_{1}$ is not an absolutely algebraic field, we deduce that $E_{1}$ is an infinite ring. Consequently $\mathbb{I}(M_{1})/M_{1}$ is an infinite ring. Thus by Theorem 3.1, we obtain that $M_{1}]$ and a fortiori $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ are infinite. It follows that from Proposition 3.10(1)(1), $T$ has infinitely many maximal subrings, as desired.
$(\bf 2)$ We have two cases, if $M_{1}$ or $M_{2}$ is not an ideal of $T$. Then analogous to last part of the proof of $(1)$, we infer that $T$ has infinitely many maximal subrings. Thus suppose that $M_{1}$ and $M_{2}$ are ideals of $T$. In this case, we have the following ring isomorphisms
@@
Consequently $T/(M_{1}\cap M_{2})\cong E_{1}^{op}\times E_{1}^{op}$, as rings (note $M_{1}\neq M_{2}$, because $S_{1}\ncong S_{2}$). Since $E_{1}$ and $E_{2}$ are infinite division rings, we deduce that $T/(M_{1}\cap M_{2})$ has infinitely many maximal subrings, by Theorem 3.6. Therefore $T$ has infinitely many maximal subrings, which complete the proof. ∎
@@ 4. Rings integral over their centers
In this section we investigate on the number of maximal subrings in a ring $T$ that is integral over its center. First we need some preliminary observations and notation. Let $T$ be a ring. Then it is clear that $\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)=\operatorname{{Max}_{r}}(T)\cap\operatorname{Max}(T)=\operatorname{{Max}_{r}}(T)\cap\operatorname{{Max}_{\ell}}(T)$, we denote this set by $\operatorname{Max}_{lr}(T)$. Define $J^{\prime}(T):=\bigcap_{M\in\operatorname{Max}_{lr}(T)}M$. Then $J^{\prime}(T)$ is an ideal of $T$ and $J(T)\subseteq J^{\prime}(T)$. In particular, if $T$ is either left or right quasi-duo ring, then $J(T)=J^{\prime}(T)$. Moreover, if $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)=\{M_{1},\ldots,M_{n}\}$ is finite (which occur, in particular, when $T$ has only finitely many maximal subrings, by Proposition 3.10(5)(5)), then $J(T)=J^{\prime}(T)\cap M_{1}\cap\cdots\cap M_{n}$. Note also that by Lemma 3.13, if $T$ is an infinite ring with $J^{\prime}(T)=0$, then either $T$ has infinitely many maximal subrings or $C(T)$ is integrally closed in $T$. We now state the following result.
@@
    $|\operatorname{Max}_{lr}(T)|\leq 2^{\aleph_{0}}$. For each $M\in\operatorname{Max}_{lr}(T)$, $T/M$ is an absolutely algebraic field (thus $|T/M|\leq\aleph_{0}$).
@@
    For each $M\in\operatorname{{Max}_{\ell}}^{\prime}(T)\cup\operatorname{{Max}_{r}}^{\prime}(T)$, $J^{\prime}(T)+M=T$.
@@
$(1)$ Assume that $T$ has only finitely many maximal subrings. Thus for each $M\in\operatorname{Max}_{lr}(T)$, the division ring $T/M$ has only finitely many maximal subrings. Since $T$ is integral over its center, we conclude that for each $M\in\operatorname{Max}_{lr}(T)$, the division ring $T/M$ is integral over its center. Thus by Proposition 3.1, we conclude that $T/M$ is a field. Clearly, $T/J^{\prime}(T)$ embeds in $\prod_{M\in\operatorname{Max}_{lr}(T)}T/M$, thus $T/J^{\prime}(T)$ is a commutative reduced ring and a fortiori $N(T)\subseteq J^{\prime}(T)$. Since $T/J^{\prime}(T)$ is a commutative ring with only finitely many maximal subrings, by [10, Corollary 1.9], we deduce that $T/J^{\prime}(T)$ is a Hilbert ring. By [10, Corollary 1.5], for each $M\in\operatorname{Max}_{lr}(T)$, the field $T/M$ is an absolutely algebraic field and consequently $|T/M|\leq\aleph_{0}$. Therefore we deduce that
$|\frac{T}{J^{\prime}(T)}|\leq|\prod_{M\in\operatorname{Max}_{lr}(T)}T/M|\leq 2^{2^{\aleph_{0}}}.$
$(2)$ Since $T/J^{\prime}(T)$ is a commutative ring, the first part of $(2)$ is an immediate consequence of [10, (1) of Proposition 2.8]. For the last part is evident by the proof of $(1)$.
$(3)$ Let $M$ be a maximal left ideal of $T$ which is not an ideal of $T$. Since $T/J^{\prime}(T)$ is a commutative ring, we conclude that $J^{\prime}(T)\nsubseteq M$, at once $M+J^{\prime}(T)=T$ by maximality of $M$. ∎
As noted in the proof of the previous proposition, if $R$ is a commutative ring with $|\operatorname{Max}(R)|>2^{\aleph_{0}}$, then $|\operatorname{RgMax}(R)|\geq 2^{\aleph_{0}}$, see [10, Proposition 2.8(1)]. For noncommutative rings we obtain the following analogue.
@@
We prove $(1)$ and $(3)$. The proof of $(2)$ is similar to $(1)$.
$(1)$ First note that if $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is infinite, then by Proposition 3.10(1)(1), we deduce that $T$ has infinitely many maximal subrings. Hence assume that $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ is finite. Thus $|\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)|=|\operatorname{{Max}_{\ell}}(T)|>2^{\aleph_{0}}$. Now since $\operatorname{Max}_{lr}(T)=\operatorname{{Max}_{\ell}}(T)\cap\operatorname{Max}(T)$, we deduce that $T$ has infinitely many maximal subrings by Proposition 4.1(2)(2).
$(3)$ If $\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)$ is infinite, then by Proposition 3.10(5)(5), we conclude that $T$ has infinitely many maximal subrings. Hence assume that $\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)$ is finite and therefore $|\operatorname{Max}(T)\cap\operatorname{{Max}_{\ell}}(T)|=|\operatorname{Max}(T)|>2^{\aleph_{0}}$. By an analogous proof of $(1)$, we obtain that $T$ has infinitely many maximal subrings. ∎
As the following main result shows, if a ring $T$ is an algebraic $K$-algebra (and therefore is integral over its center), then either $T$ has infinitely many maximal subrings, or $T$ exhibits algebraic properties similar to commutative rings with only finitely many maximal subrings, see [10].
@@
    $T$ is a quasi-duo ring. In particular, $J(T)=\bigcap_{M\in\operatorname{Max}(T)}M$.
@@
    For each $M\in\operatorname{Max}(T)$, the ring $T/M$ is an absolutely algebraic field which contains a copy of $K$. In particular, $K$ is an absolutely algebraic field and $|K|=|T/M|=\aleph_{0}$. Moreover, if $D$ is a division ring which is a subring of $T$, then $D$ is an absolutely algebraic field.
@@
    $T/J(T)$ is a reduced Hilbert commutative ring. In fact, $T/J(T)$ is a von Neumann regular ring which is integral over its prime subring (i.e., $\mathbb{Z}_{p}$, where $p=\operatorname{Char}(K)$ is a prime number). Moreover, $T$ is integral over $\mathbb{Z}_{p}$.